if you cannot find a solution label the lines as parallel or coinciding and explain why y=3x-9 2x-y=6
Answer by rfer(16322)

2x-3x-9=6
-x=15
x=-15
-------------
2*-15-y=6
-30-y=6
-y=36
y=-36
-------------
y=3*-15-9
y=-45-9
y=-54
-----------
no solution